# Completely entangled subspaces of entanglement depth $k$

### v2

@@ Abstract
We introduce a new class of entangled subspaces—completely entangled subspaces of entanglement depth $k$ ($k$-CESs). These are subspaces of multipartite Hilbert spaces containing only pure states with the entanglement depth at least $k$. We present an efficient construction of $k$-CESs of any achievable dimensionality in any multipartite scenario. Further, we discuss the relation between these subspaces and unextendible product bases (UPBs). In particular, we establish that there is a non-trivial bound on the cardinality of a UPB whose orthocomplement is a $k$-CES. Further, we discuss existence of such UPBs for qubit systems.
@@ I Introduction
Entanglement has long been known to be an enabling resource for a variety of information processing protocols such as for instance quantum communication [1] or quantum metrology [2, 3]. The ongoing rapid development of quantum network technologies requires a thorough theoretical characterization of its multipartite facet. It is a very challenging problem as the description of quantum systems with multiple nodes is far more complex than that of systems with only two subsystems [4, 5]. The topic has attracted much attention over the recent years and many profound results touching its different aspects have been reported in the literature (see, e.g., [6, 7, 8, 9, 10, 11]). Noteworthy, the theoretical advancements go largely in parallel with a remarkable progress in the experimental domain (see, e.g., [12, 13, 14, 15, 16]). Still, despite all the efforts, many aspects of many body entanglement remain insufficiently explored and further research towards its better understanding is highly demanded. This work is in line with this important trend.
@@
A particular useful tool for the characterization of entanglement in multipartite systems is that of the entanglement depth [17], which tells one how many particles in a given ensemble are genuinely entangled. The concept finds applications for example in the domain of cold gases [18, 19, 20]. The notion of entanglement can also be meaningfully considered for subspaces, not only for individual states. The most general notion is that of completely entangled subspaces (CESs), which are those subspaces that contain only pure entangled states [21, 22, 23, 24]. More specific classes of CESs considered in the literature are, e.g., subspaces containing only states with bounded tensor rank ($r$-entangled subspaces) [25, 26], or genuinely entangled subspaces [22, 29], which are particularly interesting in the multipartite scenario as they contain only genuinely entangled states—the most resourceful states in this framework. Interesting examples of genuinely entangled subspaces are for instance those spanned by the stabilizer error correction codes, including the five-qubit [27] or the toric code [28].
@@
Entangled subspaces play an important role in quantum information science making them objects of intrinsic interest (see, e.g., [30, 31, 32], for some recent results). First of all, they constitute an invaluable tool in the theory of entanglement as they can be readily utilized for a construction of mixed entangled states owing to the simple fact that any state supported on an entangled subspace is necessarily entangled. An important class of such constructed states comprises of those built from unextendible product bases [33] and having positive partial transpose (PPT). These state share an appealing property of being bound entangled, i.e., their entanglement cannot be distilled into singlets with local operations and classical communication (LOCC). In fact, this was the first general construction of multipartite bound entangled states harnessing the notion of completely entangled subspaces. Second, it was shown that almost all subspaces are completely entangled provided they are not too large, i.e., their dimension does not exceed the maximal permissible one for CESs [24]. This result was utilized to prove that almost all sets of states (separable or entangled) are locally unambiguously distinguishable if the number of elements in the set is not larger that the difference between the dimension of the whole system and the maximal dimension of a CES in the given setup [34, 24]. Moreover, entangled subspaces, such as the antisymmetric one or that introduced in Ref. [22], have been exploited to provide counterexamples to the additivity of the minimum output Rényi entropy of quantum channels [35, 36]. Their practical applicability adds further to their significance. It has been recognized that they are relevant to the dynamically growing field of quantum error correction [37, 38]. A recent development in the research exploring this connection revealed that there exists a certain trade-off relating larger capability of a code to correct errors to higher entanglement of the code space [39]. Furthermore, a link between entangled subspaces and self-testing has been established [40, 41, 42, 43] pointing to their use, e.g., in cryptography.
@@
Our current contribution is to marry the notions of the entanglement depth and entangled subspaces. Precisely, we introduce completely entangled subspaces of entanglement depth $k$ ($k$-CESs), as those containing only pure states whose entanglement depth is at least $k$. We find the maximal dimensions of such subspaces and present their universal construction from sets of non-orthogonal product vectors in any multipartite setup. Further, we discuss the relation between $k$-CESs and unextendible product bases (UPBs). This includes a derivation of a non-trivial bound on the cardinalities of UPBs leading to $k$-CESs and a discussion with a positive conclusion about their existence.
@@ II Preliminaries
We will mainly focus on the scenario with equal local dimensions $d_{i}=d$ in which case the space will be denoted as $\mathcal{H}_{d^{n}}$. Single-partite subsystems (parties) are denoted $_{1},A_{2},\dots,A_{n}:=\textbf{A}$ A. A division of the set of parties into $K$ non-overlapping non-empty sets $S_{i}$ such that $\bigcup_{i}S_{i}=\textbf{A}$ is called a $K$-partition and denoted $S_{1}|\dots|S_{K}$.
@@
where $|\phi_{i}\rangle$’s are at most $k$-partite states. If a $k$-producible state is not $(k-1)$-producible at the same time it is said to have the entanglement depth $k$ [17, 44]. States with the entanglement depth $k=1$ are named fully product as they are products of pure states on individual subsystems $A_{i}$. All other states with $k\geq 2$ are said to be entangled and those with the entanglement depth equal $n$, in particular, are called genuinely multipartite entangled (GME). These are the states which cannot be written as a product for any bipartition of the parties. Excellent examples of GME states are the Greenberger-Horne-Zeilinger (GHZ) states [45],
@@
that belong to a general class of GME states called the graph states [46].
@@
Entangled subspaces are closely related to so-called unextendible product bases and we will extensively exploit this connection in further parts of the paper. Below we remind the basic necessary facts about the discussed notions.
@@
where |±$|\pm\rangle=(|0\rangle\pm|1\rangle)/\sqrt{2}$. One can easily verify that $\mathcal{S}$ is indeed a UPB and only entangled vectors belong to $(\mathrm{span}\>\mathcal{S})^{\perp}$, which itself is thus a CES. Notably, the latter subspace is not a GES as it contains biproduct vectors, e.g., $|1\rangle\otimes|\phi\rangle$, where $|\phi\rangle$ is an entangled two-qubit vector orthogonal to |+|+$\rangle$| + - ⟩, $|1+\rangle$, and $|-1\rangle$.
Concluding this section, we give two results regarding unextendibility of a set of product vectors that will be our main tools in further parts. The first of them is the following observation.
@@
This elegant result often serves as a basis for constructions of UPBs and it will be used by us for a direct check if one of the considered bases possesses certain property of unextendibility. Its immediate consequence is a lower bound on the cardinalities of unextendible sets:
@@
Importantly, both results regarding unextendibility, Fact 1 and Corollary 1, are also applicable to sets of non-orthogonal vectors.
@@ III Completely entangled subspaces of entanglement depth $k$ ($k$-CES)
Let us remark that for $k=2$ the above definition recovers that of completely entangled subspaces, whereas for $k=n$ that of genuinely entangled subspaces.
@@
As the first example let us take the discussed above CES complementary to the UPB given in Eq. (5). Within our new terminology we can now say that it is a $2$-CES (CES), but it is not a $3$-CES (GES) because, to recall, the complementary subspace contains a three-qubit vector with the entanglement depth two.
@@
where $|\overline{\mathrm{GHZ}}_{3}\rangle=(1/\sqrt{2})(|000\rangle-|111\rangle)$ and $|\mathrm{GHZ}_{3}\rangle$ is defined in Eq. (3). Now, the subspace $\mathrm{span}\{|\psi_{0}\rangle,|\psi_{1}\rangle\}$ is a $3$-CES because the entanglement depth of both $|\psi_{0}\rangle$ and $|\psi_{1}\rangle$ is three, and, at the same time, any linear combination of these vectors is GME, i.e., it has the entanglement depth equal to nine.
As the final example we give an elementary general construction. With this aim take a GES of a $k$-partite Hilbert space. Now, take a tensor product of all the vectors from this subspace with an arbitrary, possibly different for each vector, fully product $(n-k)$-partite vector. This clearly results in a $k$-CES of an $n$-partite Hilbert space. This rather trivial method does not lead to $k$-CESs of the maximal dimensions for $k<n$ as can be seen from dimension consideration (see Appendix B) and in Section IV we will provide a more elaborate construction which does this job.
One of the central problems of the theory of entangled subspaces is the problem of verification whether a given subspace is entangled and further determining to which class it belongs to. Noteworthy, some tools have been developed in the literature so far, which could also be applied for the case of $k$-CESs. Precisely, the simple criterion based on the entanglement of basis vectors from Ref. [49] is directly applicable and the hierarchical method from Ref. [31] can be readily adapted. Another vital problem in the area is the quantification of subspace entanglement (see [50] for recent results). Here, also, we find certain methods which can be used in the current case, e.g., the SDP bounds of Ref. [51].
Let us now investigate how big—in terms of their dimensions—$k$-CESs can be. The theory of completely entangled subspaces helps us address this question successfully. For simplicity, we concentrate on the case of equal local dimensions in our derivation but a properly adjusted argument is applicable for the general case too. We have the following proposition.
@@
It was shown in Refs. [21, 22] that the maximal attainable dimension of a CES of $\mathcal{H}_{d_{1}\dots d_{n}}$ is given by $\prod_{i=1}^{n}d_{i}-[\sum_{i=1}^{n}(d_{i}-1)+1$] (notice that the term in the square bracket happens to match the RHS of Eq. (6)). Applying this result to (11), we obtain
@@
which must further be minimized (equivalently, the term in the square brackets must be maximized) over $n_{i}$’s to obtain an upper bound on the maximal dimension of a $k$-CES. This is achieved for $n_{i}=k-1$, $i=1,2,\dots,t$, $n_{t+1}=n-t(k-1)$, where $t=\left\lfloor n/(k-1)\right\rfloor$ (see Appendix A). The minimal value of (12) is thus
@@ IV Universal construction of $k$-CESs
While random subspaces do have significance in quantum information theory (see, e.g., Ref. [52]), it is the explicit constructions, which often help one analyze these objects in more detail, for example, compute entanglement measures of states supported on them. In this section, we provide a universal analytical construction of $k$-CESs, that is, we introduce a method of building $k$-CESs of any dimension for any combination of the numbers $k$, $n$, and $d$. With this aim we utilize the approach put forward recently in [56], which was originally designed for a construction of GESs. It turns out that it is versatile enough to be translated directly to the current case.
@@
Similarly to [56] one could also use DFT matrices in the construction. The method presented here can be viewed as a construction of ”non-orthgonal” $\mathrm{UPB}_{k-1}$’s (see Section V), whose elements are not mutually orthogonal.
Observe that the construction is flexible in that for a given $k$-CES we can always increase $k$ by adding some number of Vandermonde vectors with properly chosen nodes to the construction set (21). For example, consider an eleven-dimensional $2$-CES, or, simply, a CES in the standard terminology, in the system of four qubits. Its construction requires five Vandermonde vectors. If we now add two more vectors to (21) we will obtain a nine-dimensional 3-CES (in Appendix C we give an example of such subspace). If we add again two vectors we will obtain a seven-dimensional 4-CES, i.e. a GES, of the maximal permissible dimension. Further enlargement of the set of Vandermonde vectors leads to GESs of smaller dimensions.
@@ V $k$-CES and unextendible product bases
We are interested in a certain type of UPBs, namely those bases which are unextendible with vectors of the entanglement depth less than $k$ and as such leading to $k$-CESs. We thus propose the following.
@@
A UPB which is unextendible with $(k-1)$-producible vectors, i.e., states with the entanglement depth less than or equal to $k-1$, is called a $(k-1)$-unextendible product basis, denoted $\mathrm{UPB}_{\mathrm{k-1}}$.
For $k=2$ we have the standard UPBs (see Section II), while for $k=n$—genuinely unextendible product bases (GUPBs), that is UPBs unextendible even with biproduct vectors [29, 57], whose existence is currently unknown (see Refs. [57, 58, 59] for recent results in this direction).
By definition, the orthocomplement of a $\mathrm{UPB}_{\mathrm{k-1}}$ is a $k$-CES. Importantly, the reverse statement is not necessarily true even in the cases when the dimension of a $k$-CES is such that its complement may admit a $\mathrm{UPB}_{\mathrm{k-1}}$. In fact, the orhtocomplement of a $k$-CES may also be a $k$-CES itself.
@@
$\small m^{theor.}_{min}=\begin{cases}d^{n}&d=2\;\mathrm{and\;}k=n\\ t(d^{k-1}-1)+d^{n-(k-1)t}&(\mathrm{odd}\;d)\;\mathrm{or}\;[\mathrm{even\;}d\;\mathrm{and}\;\{(k-1|n\;\mathrm{and\;odd}\;t)\;\mathrm{or}\;(k-1\nmid n\;\mathrm{and\;even\;}t)\}]\\ t(d^{k-1}-1)+d^{n-(k-1)t}+1&\mathrm{otherwise}\end{cases}.$ (25)

### v3

@@ Abstract
We introduce a class of entangled subspaces: completely entangled subspaces of entanglement depth $k$ ($k$-CESs). These are subspaces of multipartite Hilbert spaces containing only pure states with an entanglement depth of at least $k$. We present an efficient construction of $k$-CESs of any achievable dimensionality in any multipartite scenario. Further, we discuss the relation between these subspaces and unextendible product bases (UPBs). In particular, we establish that there is a non-trivial bound on the cardinality of a UPB whose orthocomplement is a $k$-CES. Further, we discuss the existence of such UPBs for qubit systems.
@@ I Introduction
Entanglement has long been known to be an enabling resource for a variety of information processing protocols such as for instance quantum communication [1] or quantum metrology [2, 3]. The ongoing rapid development of quantum network technologies requires a thorough theoretical characterization of its multipartite facet. It is a very challenging problem as the description of quantum systems with multiple nodes is far more complex than that of systems with only two subsystems [4, 5]. The topic has attracted much attention over the recent years and many profound results touching its different aspects have been reported in the literature (see, e.g., [6, 7, 8, 9, 10, 11]). Noteworthy, the theoretical advancements go largely in parallel with a remarkable progress in the experimental domain (see, e.g., [12, 13, 14, 15, 16]). Still, despite all the efforts, many aspects of many-body entanglement remain insufficiently explored and further research to better understand it is highly necessary. This work is in line with this important trend.
@@
A particular useful tool for the characterization of entanglement in multipartite systems is the entanglement depth [17], which indicates how many particles in a given ensemble are genuinely entangled. The concept has found applications for example in the domain of cold gases [18, 19, 20]. The notion of entanglement can also be meaningfully considered for subspaces in addition to individual states. The most general notion is that of completely entangled subspaces (CESs), which are those subspaces that contain only pure entangled states [21, 22, 23, 24]. More specific classes of CESs considered in the literature are, e.g., subspaces containing only states with bounded tensor rank ($r$-entangled subspaces) [25, 26], or genuinely entangled subspaces [22, 29], which are particularly interesting in the multipartite scenario as they contain only genuinely entangled states—the most resourceful states in this framework. Interesting examples of genuinely entangled subspaces are for instance those spanned by the stabilizer error correction codes, including the five-qubit [27] and toric [28] codes.
@@
Entangled subspaces play an important role in quantum information science making them objects of intrinsic interest (see, e.g., [30, 31, 32], for some recent results). First of all, they constitute an invaluable tool in the theory of entanglement because they can be readily utilized for a construction of mixed entangled states owing to the simple fact that any state supported on an entangled subspace is necessarily entangled. An important class of such constructed states comprises those built from unextendible product bases [33] and having positive partial transpose (PPT). These states share the appealing property of being bound entangled, i.e., their entanglement cannot be distilled into singlets with local operations and classical communication (LOCC). In fact, this was the first general construction of multipartite bound entangled states harnessing the notion of completely entangled subspaces. Second, it was shown that almost all subspaces are completely entangled provided they are not too large, i.e., their dimension does not exceed the maximal permissible one for CESs [24]. This result was utilized to prove that almost all sets of states (separable or entangled) are locally unambiguously distinguishable if the number of elements in the set is not larger than the difference between the dimension of the whole system and the maximal dimension of a CES in the given setup [34, 24]. Moreover, entangled subspaces, such as the antisymmetric one and that introduced in Ref. [22], have been exploited to provide counterexamples to the additivity of the minimum output Rényi entropy of quantum channels [35, 36]. Their practical applicability further adds to their significance. It has been recognized that they are relevant in the dynamically growing field of quantum error correction [37, 38]. A recent development in the research exploring this connection revealed that there exists a certain trade-off relating the larger capability of a code to correct errors to higher entanglement of the code space [39]. Furthermore, a link between entangled subspaces and self-testing was established [40, 41, 42, 43] pointing to their use, e.g., in cryptography.
@@
Our current contribution is to marry the notions of the entanglement depth and entangled subspaces. Precisely, we introduce completely entangled subspaces of entanglement depth $k$ ($k$-CESs), as those subspaces containing only pure states whose entanglement depth is at least $k$. We find the maximal dimensions of such subspaces and present their universal construction from sets of non-orthogonal product vectors in any multipartite setup. Further, we discuss the relation between $k$-CESs and unextendible product bases (UPBs). This includes the derivation of a non-trivial bound on the cardinalities of UPBs leading to $k$-CESs and a discussion with a positive conclusion about their existence.
@@ II Preliminaries
We will mainly focus on the scenario with equal local dimensions $d_{i}=d$ in which case the space will be denoted as $\mathcal{H}_{d^{n}}$. Single-partite subsystems (parties) are denoted $_{1},A_{2},\dots,A_{n}:=\textbf{A}$ A. A division of the set of parties into $K$ non-overlapping non-empty sets $S_{i}$ such that $\bigcup_{i}S_{i}=\textbf{A}$ is called a $K$-partition and is denoted $S_{1}|\dots|S_{K}$.
@@
where $|\phi_{i}\rangle$’s are at most $k$-partite states. If a $k$-producible state is not $(k-1)$-producible at the same time it is said to have entanglement depth $k$ [17, 44]. States with entanglement depth $k=1$ are named fully product because they are products of pure states on individual subsystems $A_{i}$. All other states with $k\geq 2$ are said to be entangled and those with entanglement depth equal $n$, in particular, are called genuinely multipartite entangled (GME). These are the states which cannot be written as a product for any bipartition of the parties. Excellent examples of GME states are the Greenberger-Horne-Zeilinger (GHZ) states [45],
@@
which belong to a general class of GME states called graph states [46].
@@
Entangled subspaces are closely related to the so-called unextendible product bases and we will extensively exploit this connection in later parts of the paper. Below we remind the basic necessary facts about the discussed notions.
@@
where |±$|\pm\rangle=(|0\rangle\pm|1\rangle)/\sqrt{2}$. One can easily verify that $\mathcal{S}$ is indeed a UPB and only entangled vectors belong to $(\mathrm{span}\>\mathcal{S})^{\perp}$, which itself is thus a CES. Notably, the latter subspace is not a GES because it contains biproduct vectors, e.g., $|1\rangle\otimes|\phi\rangle$, where $|\phi\rangle$ is an entangled two-qubit vector orthogonal to |+|+$\rangle$| + - ⟩, $|1+\rangle$, and $|-1\rangle$.
Concluding this section, we give two results regarding the unextendibility of a set of product vectors that will be our main tools in later parts. The first is the following observation.
@@
This elegant result often serves as a basis for the constructions of UPBs and it will be used by us for a direct check of whether one of the considered bases possesses certain property of unextendibility. Its immediate consequence is a lower bound on the cardinalities of unextendible sets:
@@
Importantly, both results related to unextendibility, Fact 1 and Corollary 1, are also applicable to sets of non-orthogonal vectors.
@@ III Completely entangled subspaces of entanglement depth $k$ ($k$-CES)
Let us remark that for $k=2$ the above definition recovers the definition of completely entangled subspaces, whereas for $k=n$ it recovers the definition of genuinely entangled subspaces.
@@
As the first example let us take the above-discussed CES complementary to the UPB given in Eq. (5). Within our terminology we can now say that it is a $2$-CES (CES), but it is not a $3$-CES (GES) because, as discussed, the complementary subspace contains a three-qubit vector with an entanglement depth of $2$.
@@
where $|\overline{\mathrm{GHZ}}_{3}\rangle=(1/\sqrt{2})(|000\rangle-|111\rangle)$ and $|\mathrm{GHZ}_{3}\rangle$ is defined in Eq. (3). Now, the subspace $\mathrm{span}\{|\psi_{0}\rangle,|\psi_{1}\rangle\}$ is a $3$-CES because the entanglement depth of both $|\psi_{0}\rangle$ and $|\psi_{1}\rangle$ is $3$, and, at the same time, any linear combination of these vectors is GME, i.e., it has an entanglement depth equal to $9$.
As the final example we give an elementary general construction. With this aim take a GES of a $k$-partite Hilbert space. Now, take a tensor product of all the vectors from this subspace with an arbitrary, possibly different for each vector, fully product $(n-k)$-partite vector. This clearly results in a $k$-CES of an $n$-partite Hilbert space. This rather trivial method does not lead to $k$-CESs of the maximal dimensions for $k<n$ as can be seen from dimension consideration (see Appendix B) and in Section IV we will provide a more elaborate construction which does this job. However, it is noteworthy that if we choose the otherwise arbitrary vectors to be the same for each of the vectors from a GES, the resulting subspace will be a $k$-CES with all vectors belonging to it having an entanglement depth equal exactly to $k$.
One of the central problems of the theory of entangled subspaces is the problem of verifying whether a given subspace is entangled and further determining to which class it belongs. It is worth noting that some tools have been developed in the literature so far, which could also be applied for the case of $k$-CESs. Specifically, the simple criterion based on the entanglement of basis vectors from Ref. [49] is directly applicable and the hierarchical method from Ref. [31] can be readily adapted. Another vital problem in the area is the quantification of subspace entanglement (see [50] for recent results). Here, also, we find certain methods which can be used in the current case, e.g., the semidefinite programming bounds of Ref. [51].
Let us now investigate how big—in terms of their dimensions—$k$-CESs can be. The theory of completely entangled subspaces helps us address this question successfully. For simplicity, we concentrate on the case of equal local dimensions in our derivation but a properly adjusted argument is applicable in the general case too. We have the following proposition.
@@
It was shown in Refs. [21, 22] that the maximal attainable dimension of a CES of $\mathcal{H}_{d_{1}\dots d_{n}}$ is given by $\prod_{i=1}^{n}d_{i}-[\sum_{i=1}^{n}(d_{i}-1)+1$] (notice that the term in the square bracket happens to match the right-hand side of Eq. (6)). Applying this result to (11), we obtain
@@
which must further be minimized (equivalently, the term in the square brackets must be maximized) over $n_{i}$’s to obtain an upper bound on the maximal dimension of a $k$-CES. This is achieved for $n_{i}=k-1$, $i=1,2,\dots,t$, and $n_{t+1}=n-t(k-1)$, where $t=\left\lfloor n/(k-1)\right\rfloor$ (see Appendix A). The minimal value of (12) is thus
@@ IV Universal construction of $k$-CESs
While random subspaces do have significance in quantum information theory (see, e.g., Ref. [52]), it is the explicit constructions, which often help one analyze these objects in more detail, for example, to compute entanglement measures of states supported on them. In this section, we provide a universal analytical construction of $k$-CESs, that is, we introduce a method for building $k$-CESs of any dimension for any combination of the numbers $k$, $n$, and $d$. With this aim we utilize the approach put forward recently in [56], which was originally designed for the construction of GESs. It turns out that it is versatile enough to be translated directly to the current case.
@@
Similar to [56] one could also use discrete Fourier transform matrices in the construction. The method presented here can be viewed as a construction of ”non-orthgonal” $\mathrm{UPB}_{k-1}$’s (see Section V), whose elements are not mutually orthogonal.
Observe that the construction is flexible in that for a given $k$-CES we can always increase $k$ by adding some number of Vandermonde vectors with properly chosen nodes to the construction set (21). For example, consider an eleven-dimensional $2$-CES, or, simply, a CES in the standard terminology, in the system of four qubits. Its construction requires five Vandermonde vectors. If we now add two more vectors to (21) we will obtain a nine-dimensional 3-CES (in Appendix C we give an example of such a subspace). If we again add two vectors we will obtain a seven-dimensional 4-CES, i.e. a GES, of the maximal permissible dimension. Further enlargement of the set of Vandermonde vectors leads to GESs of smaller dimensions.
@@ V $k$-CES and unextendible product bases
We are interested in a certain type of UPBs, namely those bases which are unextendible with vectors of the entanglement depth less than $k$ that thus lead to $k$-CESs. We thus propose the following.
@@
A UPB which is unextendible with $(k-1)$-producible vectors, i.e., states with an entanglement depth less than or equal to $k-1$, is called a $(k-1)$-unextendible product basis, denoted $\mathrm{UPB}_{\mathrm{k-1}}$.
For $k=2$ we have the standard UPBs (see Section II), while for $k=n$— we have genuinely unextendible product bases (GUPBs), that is UPBs unextendible even with biproduct vectors [29, 57], whose existence is currently unknown (see Refs. [57, 58, 59] for recent results related to this topic).
By definition, the orthocomplement of a $\mathrm{UPB}_{\mathrm{k-1}}$ is a $k$-CES. Importantly, the reverse statement is not necessarily true even in the cases when the dimension of a $k$-CES is such that its complement may admit a $\mathrm{UPB}_{\mathrm{k-1}}$. In fact, the orthocomplement of a $k$-CES may also be a $k$-CES itself.
@@
$\small m^{theor.}_{min}=\begin{cases}d^{n}&d=2\;\mathrm{and\;}k=n,\\ t(d^{k-1}-1)+d^{n-(k-1)t}&(\mathrm{odd}\;d)\;\mathrm{or}\;[\mathrm{even\;}d\;\mathrm{and}\;\{(k-1|n\;\mathrm{and\;odd}\;t)\;\mathrm{or}\;(k-1\nmid n\;\mathrm{and\;even\;}t)\}],\\ t(d^{k-1}-1)+d^{n-(k-1)t}+1&\mathrm{otherwise}.\end{cases}$ (25)